OpenStudy (anonymous):

Algebra 1, please help me :) Create your own quadratic equation that has an axis of symmetry of x = −1 and whose graph opens down. Then find the vertex, domain, range, and x-intercepts. Show all work to receive full credit.

OpenStudy (anonymous):

cant understand is there any graph??

OpenStudy (anonymous):

No graph...

OpenStudy (anonymous):

Random graph. Formula for the axis of symmetry: \[x = -\frac {b}{2a}\]\[x = -1\]\[-1 = -\frac b{2a}\] \[y = -x^2 - 2x - 1\]Vertex: (-1, 0) Domain: All real numbers Range: y ≤ 0 x - intercepts: \[y = -x^2 - 2x - 1\]\[0 = -x^2 - 2x - 1\]\[0 = x^2 + 2x + 1\]\[0 = (x + 1)(x + 1)\]\[x = -1